# In-Context Semi-Supervised Learning

### v1

@@ Appendix B Derivation of the GD Update Equation for ICL with Categorical Observations
Our derivation is based on the assumption that $f(\phi)$ resides in a reproducing kernel Hilbert space (RKHS) scholkopf2002learning, but the setup extends to softmax-based attention kernels as well Aaron_ICL. From the RKHS perspective, let $f(\phi)=A\psi(\phi)$, with $\psi(\phi)$ a fixed mapping of features $\phi$ to a Hilbert space, and the parameters acting in that space are associated with matrix $A$.
@@ E.2 Product Manifolds
The five classical manifolds introduced in Section E.1 are all low‐dimensional and possess relatively simple label boundaries. In practical applications, however, data often reside on much higher‐dimensional nonlinear spaces whose intrinsic geometry cannot be recovered from any single elementary chart. By taking the Cartesian product of elementary manifolds, we can construct highly complex manifolds in arbitrary dimensions. These product manifolds play an important role in practical applications ranging from computational chemistry (zhang2021product) and image and 3D structure analysis (fumero2021learning). Evaluating on these products therefore probes the algorithmic generalisability of our Transformer—namely, whether it can still recover the manifold structure and correct labels when confronted with far higher ambient dimension and intricate decision boundaries.
@@ E.3 Image Manifold
To evaluate our approach on high-dimensional, we created image manifolds using Stable Diffusion v1.5 Rombach_2022_CVPR. These manifolds represent a significant increase in complexity and dimensionality compared to our synthetic manifolds, testing the transfer capabilities of our semi-supervised in-context learning method.
@@ F.1 Dependence on ICL Transformer Depth
The results demonstrate that the 2-layer ICL architecture consistently outperforms the 1-layer architecture across all labeled data percentages. This supports the theoretical framework in Aaron_ICL, showing that increased depth allows the model to learn more sophisticated manifold representations. The performance improvement is more pronounced for the RBF kernel, which further emphasizes the kernel’s ability to capture non-linear relationships in the data manifold. We observe that for both kernel types, the 2-layer ICL architecture provides a consistent advantage, with RBF kernels outperforming linear counterparts at both depths. This advantage is most notable in the low-label regime (5-15% labeled data), highlighting the importance of architectural capacity when labeled data is scarce.
@@ F.5 ImageNet100 Experiments with $39\%$ Labels
Figure 10: In-context learning on ImageNet100 with $39\%$ labels and dataset size 50005000. Each dataset consists of 200 batches, with each batch containing 100 samples (50 images per class) represented by VGG-29 features. The left column reports ICL accuracy, and the right column reports the separation score (intra-class minus inter-class similarity, following caron2018deep; oord2018representation). We compare three methods: Orig+E2E-ICL (red, PE from our model), a Transformer baseline (blue, PE from the final transformer layer min2022metaicl), and Eig+ICL (green, features pre-extracted from VGG-29).
@@
The separation score is defined as the difference between average intra-class similarity and inter-class similarity caron2018deep; oord2018representation:
@@ F.7 Models Under Evaluation
Kernel logistic regression on the original coordinates using the Gaussian kernel $\kappa_{\text{RBF}}(x,x^{\prime})=\exp(-\gamma\|x-x^{\prime}\|_{2}^{2})$ scholkopf2002learning. The decision function is $\sigma\!\bigl(\sum_{j=1}^{n}\alpha_{j}\kappa_{\text{RBF}}(x_{i},x_{j})\bigr)$ with an $\ell_{2}$ penalty on $\bm{\alpha}$, providing a nonlinear classifier refit at each label ratio.
@@ Appendix G Transformer Baseline Architecture
Model: SmallScaleTransformer (adapted from vonoswald2023transformers).   This serves as an in-context Transformer pipeline, where raw features and tokenized labels are jointly embedded, processed by a Transformer encoder, and predictions are made in-context.

### v2

@@ Appendix B Derivation of the GD Update Equation for ICL with Categorical Observations
Our derivation is based on the assumption that $f(\phi)$ resides in a reproducing kernel Hilbert space (RKHS) Schölkopf and Smola (2002), but the setup extends to softmax-based attention kernels as well Wang et al. (2025). From the RKHS perspective, let $f(\phi)=A\psi(\phi)$, with $\psi(\phi)$ a fixed mapping of features $\phi$ to a Hilbert space, and the parameters acting in that space are associated with matrix $A$.
@@ E.2 Product Manifolds
The five classical manifolds introduced in Section E.1 are all low‐dimensional and possess relatively simple label boundaries. In practical applications, however, data often reside on much higher‐dimensional nonlinear spaces whose intrinsic geometry cannot be recovered from any single elementary chart. By taking the Cartesian product of elementary manifolds, we can construct highly complex manifolds in arbitrary dimensions. These product manifolds play an important role in practical applications ranging from computational chemistry (Zhang et al., 2021) and image and 3D structure analysis (Fumero et al., 2021). Evaluating on these products therefore probes the algorithmic generalisability of our Transformer—namely, whether it can still recover the manifold structure and correct labels when confronted with far higher ambient dimension and intricate decision boundaries.
@@ E.3 Image Manifold
To evaluate our approach on high-dimensional, we created image manifolds using Stable Diffusion v1.5 Rombach et al. (2022). These manifolds represent a significant increase in complexity and dimensionality compared to our synthetic manifolds, testing the transfer capabilities of our semi-supervised in-context learning method.
@@ F.1 Dependence on ICL Transformer Depth
The results demonstrate that the 2-layer ICL architecture consistently outperforms the 1-layer architecture across all labeled data percentages. This supports the theoretical framework in Wang et al. (2025), showing that increased depth allows the model to learn more sophisticated manifold representations. The performance improvement is more pronounced for the RBF kernel, which further emphasizes the kernel’s ability to capture non-linear relationships in the data manifold. We observe that for both kernel types, the 2-layer ICL architecture provides a consistent advantage, with RBF kernels outperforming linear counterparts at both depths. This advantage is most notable in the low-label regime (5-15% labeled data), highlighting the importance of architectural capacity when labeled data is scarce.
@@ F.5 ImageNet100 Experiments with $39\%$ Labels
Figure 10: In-context learning on ImageNet100 with $39\%$ labels and dataset size 50005000. Each dataset consists of 200 batches, with each batch containing 100 samples (50 images per class) represented by VGG-29 features. The left column reports ICL accuracy, and the right column reports the separation score (intra-class minus inter-class similarity, following Caron et al. (2018); Oord et al. (2018)). We compare three methods: Orig+E2E-ICL (red, PE from our model), a Transformer baseline (blue, PE from the final transformer layer Min et al. (2022)), and Eig+ICL (green, features pre-extracted from VGG-29).
@@
The separation score is defined as the difference between average intra-class similarity and inter-class similarity Caron et al. (2018); Oord et al. (2018):
@@ F.7 Models Under Evaluation
Kernel logistic regression on the original coordinates using the Gaussian kernel $\kappa_{\text{RBF}}(x,x^{\prime})=\exp(-\gamma\|x-x^{\prime}\|_{2}^{2})$ Schölkopf and Smola (2002). The decision function is $\sigma\!\bigl(\sum_{j=1}^{n}\alpha_{j}\kappa_{\text{RBF}}(x_{i},x_{j})\bigr)$ with an $\ell_{2}$ penalty on $\bm{\alpha}$, providing a nonlinear classifier refit at each label ratio.
@@ Appendix G Transformer Baseline Architecture
Model: SmallScaleTransformer (adapted from von Oswald et al., 2023).   This serves as an in-context Transformer pipeline, where raw features and tokenized labels are jointly embedded, processed by a Transformer encoder, and predictions are made in-context.